OpenStudy (anonymous):

The gravitational force on the solar system is the net force due to all the matter inside our orbit. Most of that matter is concentrated near the center of the galaxy. Assume that the matter has a spherical distribution, like a giant star. What is the approximate mass of the galactic center?

OpenStudy (anonymous):

what values have we got? velocity of the solar system, mass of solar system....?

OpenStudy (anonymous):

The solar system is 25,000 light years from the center of our Milky Way galaxy. One light year is the distance light travels in one year at a speed of 3.0 10^8 m/s. Astronomers have determined that the solar system is orbiting the center of the galaxy at a speed of 230 km/s.

OpenStudy (anonymous):

Sorry I didn't post that earlier, whoops!

OpenStudy (anonymous):

good. so the solar system (a point mass) is performing (assume like anything) a uniform circular motion about the galactic center (another point mass). then, the centripetal force of the solar system is balanced by the gravitational force

OpenStudy (anonymous):

\[ {\cancel{m_{\rm solar}}v^2\over \cancel{r}}=G{m_{\rm core}\cancel{m_{\rm solar}}\over r^\cancel{2}} \]

OpenStudy (anonymous):

\[v^2={Gm_{\rm core}\over r}\]

OpenStudy (anonymous):

make sure all the values are in kg, m, s, N, ...

OpenStudy (anonymous):

what answer do you come to? I got an answer and it was wrong..

OpenStudy (anonymous):

what was your answer?

OpenStudy (anonymous):

1.88*10^32

OpenStudy (anonymous):

\[\left(230\times10^3\right)^2=\frac{6.67\times10^{-11}\times m_{\rm core}}{25\times10^3\times3\times10^8}\]

OpenStudy (anonymous):

I didn't change the 230 to m/s, stupid! sorry!

OpenStudy (anonymous):

thought so, and told you so :)

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

what did you multiply the 25000 light years by?

OpenStudy (anonymous):

speed of light

OpenStudy (anonymous):

does that give distance in meters?

OpenStudy (anonymous):

I used the exact formula provided and still wrong, came to 5.9*10^33

OpenStudy (anonymous):

answer was 1.87*10^41

OpenStudy (anonymous):

no idea how to get it, I am a biology person, not physics haha

OpenStudy (anonymous):

my bad \[ \left(230\times10^3\right)^2=\frac{6.67\times10^{-11}\times m_{\rm core}}{25\times10^3\times3\times10^8\color{red}{\times365\times24\times3600}}=1.876\times10^{41}{\rm kg} \] http://wolfr.am/XZbB3r